OpenStudy (anonymous):

Only 2 more :p Solve y – 3x = 5 by the substitution method. y + x = 3

OpenStudy (anonymous):

Quick

OpenStudy (bahrom7893):

y = 3-x 3-4x=5 x = -1/2 y = 3/2

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

answer y = 3-x 3-4x=5 x = -1/2 y = 3/2

OpenStudy (anonymous):

AYE I POSTED SAME QUESTION 4 FN TIMES SOMEONE HELP :p pls and thanks

OpenStudy (lgbasallote):

ARGHHHHHHHH!!!!! HAHAHHA

OpenStudy (anonymous):

DANGG!

OpenStudy (mattfeury):

I think in the end it would benefit you to actually try some of these rather than just get the answers. You'll only be doing yourself a disservice by being lazy. Just my $.02.